# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

### v1

@@ Discussion.
On another note, our ordering of the Coulomb interaction (Eq. 1) is consistent with the ’average’ subtraction scheme in MATBG literature [75, 74]. Another widely used protocol is the ’graphene’ subtraction scheme [6, 9], where $\mathcal{H}_{\text{int}}$ is normal-ordered [90] with respect to the ground state of two decoupled layers at neutrality.
It has been assumed that due to the subtraction, the normal state is close to the bare ground state in the graphene scheme [6, 12, 11, 13, 89, 91]. Contrary to this expectation, our calculations with graphene subtraction, found in the Supplementary Materials [82], show significant band widening also; the flat bands reach a bandwidth of $43.8$ meV for $\epsilon_{r}=10$. More notably, we have identified a topological phase transition on the flat bands at some value of $\epsilon_{r}$ between $50$ and $10$.
The results presented here promise to have extensive consequences in our understanding of MATBG. We already mentioned the pattern of explicit symmetry breaking, $U(4)\times U(4)\to U(4)_{\text{chiral}}$ instead of $U(4)\times U(4)\to U(4)_{\text{flat}}$ which is conventionally assumed [68, 33, 32, 15]. In close connection to this, the VP and TIVC states are ’dynamically enforced’ to be gapless in contrast to most theory results [68, 92, 15].
In another regard, let us consider the strain-stabilized, moiré translations-breaking, inconmensurate Kekulé spiral (IKS) phase [86, 89, 14, 76]. The IKS momentum, $\bm{q}_{\text{IKS}}$, is determined by a nesting condition in the band structure of the nematic semimetal (NSM), and it can be measured in STM experiments [93]. The NSM can be significantly modified by the band renormalization, for instance the high energy cost if the Dirac nodes lie close to $\Gamma_{M}$ from $\mathcal{H}_{0}$ competes with the exchange gain if the Dirac cones lie at the BZ center [10]. Thus, a comparison between the theoretically predicted and the experimentally determined $\bm{q}_{\text{IKS}}$ provides a direct test of our findings.
Acknowledgements.
This work was supported by grant PID2020–113164 GBI00 funded by MCIN/AEI/10.13039/501100011033, PID2023-146461NB-I00 and PRE2021-097070 funded by Ministerio de Ciencia, Innovación y Universidades, and the CSIC Research Platform on Quantum Technologies PTI-001. The access to computational resources of CESGA (Centro de Supercomputación de Galicia) is also gratefully acknowledged.
@@ Acknowledgements.
Supplementary Material for ”Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene”
I Lattice geometry and tight-binding model
@@
We set up the twisted bilayer geometry starting from a graphene bilayer with the two layers stacked on top of each other, at vertical positions $z=\pm d_{0}/2=\pm 1.6755$. The top layer gets rotated by an angle $\theta/2$ and the bottom layer by $-\theta/2$, with the center of rotation being the center of one of the carbon hexagons. For a generic twist angle $\theta$, the resulting structure is not strictly periodic, only for particular angles there is a commensurate crystalline structure. A subset of such commensurate angles can be parameterized as $\cos(\theta)=1-1/(6n^{2}+6n+2)$ with integer $n$ [94]. The unit vectors of the resulting superlattice are $\bm{L}_{1}=L_{M}(\sqrt{3}/2,1/2)$ and $\bm{L}_{2}=L_{M}(-\sqrt{3}/2,1/2)$ with $L_{M}=a_{0}\sqrt{3n^{2}+3n+1}$ the moiré lattice constant. The reciprocal lattice in turn has unit vectors $\bm{g}_{1}=2\pi/\sqrt{3}L_{M}(1/2,\sqrt{3}/2)$ and $\bm{g}_{2}=2\pi/\sqrt{3}L_{M}(1/2,\sqrt{3}/2)$.
The magic-angle of twisted bilayer graphene sits between $1.0^{\circ}$ and $1.1^{\circ}$ in different experiments and theoretical models. In this work we chose $\theta=1.05^{\circ}$ ($n=31$) such that the bandwidth of the flat bands of our tight-binding model is minimal. The moiré lattice constant is of $13.4$ nm and the number of atoms in the unit cell is $11908$. We have included the effects of in-plane lattice relaxation, i.e. symmetry-preserving intrinsic strain, using the model of Ref. [77].
@@
$\begin{gathered}t(\bm{r})=-V_{pp\pi}(r)\big{(}1-\cos^{2}(\varphi_{\bm{r}})\big{)}+V_{pp\sigma}(r)\cos^{2}(\varphi_{\bm{r}}),\\ V_{pp\pi}(r)=V_{pp\pi}^{0}\ \text{exp}\big{(}-(r-a_{cc})/r_{0}\big{)},\quad\quad\quad V_{pp\sigma}(r)=V_{pp\sigma}^{0}\ \text{exp}\big{(}-(r-d_{0})/r_{0}\big{)},\end{gathered}$ (S1)
@@
Figure S1: (a) Top view of MATBG with a twist angle of $6.01^{\circ}$ ($n=5$). The lattice vectors $\bm{L}_{1}$, $\bm{L}_{2}$ and the moiré Wigner-Seitz cell are marked. (b) Triangular loops used for computing the valley charge. (c) Brillouin zones of the rotated top (blue) and bottom (orange) monolayers, and the hexagonal Brillouin zones of the coupled layers (black), drawn in an extended scheme. The $K(K^{\prime})$ points of the original graphenes are depicted. (d) Moiré Brillouin zone of MATBG, with the high-symmetry $\Gamma_{M}$, $M_{M}$, $K_{M}$, $K^{\prime}_{M}$ points and the reciprocal lattice vectors $\bm{g}_{1}$, $\bm{g}_{2}$ labeled.
II Mean-field decoupling and many-body projection
@@
$\begin{split}&{H}=H_{\text{TB}}+{H}_{\text{int}}\\ &H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\\ &{H}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}-\frac{1}{2}\bigg{)}+\frac{1}{2}\sum_{\bm{r}s}U\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}-\frac{1}{2}\bigg{)},\end{split}$ (S2)
@@
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\frac{1}{2}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (S3)
@@
$c^{\dagger}_{\bm{k}\bm{\delta}s}=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{R}_{\ell}}e^{i\bm{k}\cdot(\bm{\delta}+\bm{R}_{\ell})}c^{\dagger}_{\bm{\delta}+\bm{R}_{\ell},s},$ (S4)
@@
$\begin{split}{H}_{\text{MF}}=&\sum_{\bm{k}ns}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\text{constant},\\ c^{\dagger}_{\bm{k}ns}=&\sum_{\bm{\delta}}\langle\bm{k}\bm{\delta}|\bm{k}n\rangle c^{\dagger}_{\bm{k}\bm{\delta}s}=\sum_{\bm{r}}\langle\bm{r}|\bm{k}n\rangle c^{\dagger}_{\bm{r}s},\end{split}$ (S5)
where $\langle\bm{k}\bm{\delta}|\bm{k}n\rangle$ and $\langle\bm{r}|\bm{k}n\rangle$ are the wave function amplitudes (independent of spin) of mode $c^{\dagger}_{\bm{k}ns}$ in the Bloch and real-space basis, respectively. The conduction bands have band index $n=1,2,...$ with increasing $n$ corresponding to increasing energy, and $n=-1,-2,...$ denotes the valence bands with increasing $|n|$ corresponding to decreasing energy. In the self-consistent algorithm we used the ground state of the bare system ($H_{\text{TB}}$) as initial seed and imposed the crystallographic symmetries and spinless time-reversal symmetry at each iteration. We used a $6\times 6$ discretization of the Brillouin zone and the ODA algorithm [96] to accelerate convergence.
@@
$\begin{split}\mathcal{P}=&\sum_{I}|\Psi_{I}\rangle\big{\langle}\Psi_{I}|,\\ |\Psi_{I}\rangle=&\prod_{\bm{k}ns\in I}c^{\dagger}_{\bm{k}ns}\prod_{\bm{k}s,n<{-n_{B}}/{2}}c^{\dagger}_{\bm{k}ns}|0\rangle,\end{split}$ (S6)
@@
$+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})c^{\dagger}_{\alpha}c^{\dagger}_{\beta}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}+\text{others},$ (S7)
@@
$\mathcal{P}V\mathcal{P}=\frac{1}{2}\sum_{\alpha\beta\gamma\delta}V_{\alpha\beta\gamma\delta}c^{\dagger}_{\alpha}c^{\dagger}_{\beta}c_{\gamma}c_{\delta}+\sum_{\alpha\beta}\sum_{i}(V_{\alpha ii\beta}-V_{i\alpha i\beta})c^{\dagger}_{\alpha}c_{\beta}+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})$ (S8)
Using Eq. (S7) one can check that the expression in Eq. (S8) has the same action as $V$ if we restrict to initial and final states contained in $\mathcal{P}$. One now sees that, crucially, the occupied modes that have been integrated out induce new terms which are precisely the Hartree and Fock potentials acting on the central bands. The constant piece is the energy of the projected-out Fermi sea.
@@
$\mathcal{P}O\mathcal{P}=\sum_{\alpha\beta}O_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}+\sum_{i}O_{ii}.$ (S9)
@@
${\mathcal{H}}=\mathcal{P}{H}\mathcal{P}=$ $\sum_{\bm{k}ns}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\ d^{\dagger}_{\bm{r}s}d^{\dagger}_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r}s}+\sum_{\bm{r}}U\ d^{\dagger}_{\bm{r}\uparrow}d^{\dagger}_{\bm{r}\downarrow}d_{\bm{r}\downarrow}d_{\bm{r}\uparrow}$
@@
$d^{\dagger}_{\bm{r}s}={\sum_{\bm{k}n}}^{\prime}\langle\bm{k}n|\bm{r}\rangle c^{\dagger}_{\bm{k}ns},\quad\quad\quad\rho_{0\bm{r}}=\sideset{}{{}^{\prime}}{\sum}_{\bm{k},n<0}|\langle\bm{r}|\bm{k}n\rangle|^{2}.$ (S11)
@@
$\begin{split}&{\mathcal{H}}={\mathcal{H}}_{0}+{\mathcal{H}}_{\text{int}}+\text{constant},\\ &{\mathcal{H}}_{0}={\sum_{\bm{k}ns}}^{\prime}{h}(\bm{k})_{nm}c^{\dagger}_{\bm{k}ns}c_{\bm{k}ms},\\ &{\mathcal{H}}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\delta\rho_{\bm{r}s}\delta\rho_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}U\delta\rho_{\bm{r}\uparrow}\delta\rho_{\bm{r}\downarrow}.\end{split}$ (S12)
@@
$\begin{gathered}h(\bm{k})_{nm}=\varepsilon^{\text{MF}}_{n}(\bm{k})\delta_{nm}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}}}V(\bm{r}-\bm{r^{\prime}})\langle\bm{k}n|\bm{r}\rangle\langle\bm{r^{\prime}}|\bm{k}m\rangle\bigg{(}\sideset{\,}{{}^{\prime}}{\sum}_{\bm{q},\ell<0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle-\sideset{}{{}^{\prime}}{\sum}_{\bm{q},\ell>0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle\bigg{)},\\ \delta\rho_{\bm{r}s}=\rho_{\bm{r}s}-\rho_{0\bm{r}},\quad\quad\quad\quad\quad\rho_{\bm{r}s}=d^{\dagger}_{\bm{r}s}d_{\bm{r}s}.\end{gathered}$ (S13)
This is the form of the effective model studied in the main text.
III Vortex Chern basis of the flat bands

### v3

@@ V Discussion
On another note, our ordering of the Coulomb interaction (Eq. 1) is consistent with the ’average’ subtraction scheme in MATBG literature [79, 76]. Another widely used protocol is the ’graphene’ subtraction scheme [6, 9], where ${H}_{\text{int}}$ is normal-ordered [91] with respect to the ground state of two decoupled layers at neutrality.
It has been assumed that due to the subtraction, the normal state is close to the bare ground state in the graphene scheme [6, 12, 11, 13, 90, 92]. Contrary to this expectation, our calculations with graphene subtraction, found in Appendix E, show significant band widening also; the flat bands reach a bandwidth of $43.8$ meV for $\epsilon_{r}=10$. Moreover, in this scheme we have identified a topological phase transition on the flat bands for some value of $\epsilon_{r}$ between $50$ and $10$.
To conclude, let us mention that, in our model at $\theta=1.05^{\circ}$, the conduction (valence) states of the flat bands are mainly composed of conduction (valence) states of the parent graphene layers [93, 94]; the Hartree-Fock potentials, hence, drive the graphene conduction-like sates upward in energy while the valence states are lowered. This suggests that the magic angle, as defined by the angle at which there is a band inversion at $\Gamma_{M}$ [3, 94, 54, 93] (in our tight-binding model this occurs at $\theta=1.02^{\circ}$), gets renormalized towards a lower value  [^1] M. Sánchez Sánchez, T. Stauber and J. González, in preparation. The magic angle is found to be between $0.9^{\circ}$ and $1.0^{\circ}$ for $\epsilon_{r}=10$, $U=4$ eV..
Acknowledgments
We thank Patrick Ledwith and Eslam Khalaf for fruitful discussions. This work was supported by grant PID2020–113164 GBI00 funded by MCIN/AEI/10.13039/501100011033, PID2023-146461NB-I00 and PRE2021-097070 funded by Ministerio de Ciencia, Innovación y Universidades, and the CSIC Research Platform on Quantum Technologies PTI-001. The access to computational resources of CESGA (Centro de Supercomputación de Galicia) is also gratefully acknowledged.
@@ Acknowledgments
Appendix A Lattice geometry and tight-binding model
@@
We set up the twisted bilayer geometry starting from a graphene bilayer with the two layers stacked on top of each other, at vertical positions $z=\pm d_{0}/2=\pm 1.6755$. The top layer gets rotated by an angle $\theta/2$ and the bottom layer by $-\theta/2$, with the center of rotation being the center of one of the carbon hexagons. For a generic twist angle $\theta$, the resulting structure is not strictly periodic, only for particular angles there is a commensurate crystalline structure. A subset of such commensurate angles can be parameterized as $\cos(\theta)=1-1/(6n^{2}+6n+2)$ with integer $n$ [96]. The unit vectors of the resulting superlattice are $\bm{L}_{1}=L_{M}(\sqrt{3}/2,1/2)$ and $\bm{L}_{2}=L_{M}(-\sqrt{3}/2,1/2)$ with $L_{M}=a_{0}\sqrt{3n^{2}+3n+1}$ the moiré lattice constant. The reciprocal lattice in turn has unit vectors $\bm{g}_{1}=4\pi/\sqrt{3}L_{M}(1/2,\sqrt{3}/2)$ and $\bm{g}_{2}=4\pi/\sqrt{3}L_{M}(-1/2,\sqrt{3}/2)$.
The magic-angle of twisted bilayer graphene sits between $1.0^{\circ}$ and $1.1^{\circ}$ in different experiments and theoretical models. In this work we chose $\theta=1.05^{\circ}$ ($n=31$). The moiré lattice constant is of $13.4$ nm and the number of atoms in the unit cell is $11908$. We have included the effects of in-plane lattice relaxation, i.e. symmetry-preserving intrinsic strain, using the model of Ref. [77].
@@
$\begin{gathered}t(\bm{r})=-V_{pp\pi}(r)\big{(}1-\cos^{2}(\varphi_{\bm{r}})\big{)}+V_{pp\sigma}(r)\cos^{2}(\varphi_{\bm{r}}),\\ V_{pp\pi}(r)=V_{pp\pi}^{0}\ \text{exp}\big{(}-(r-a_{cc})/r_{0}\big{)},\quad\quad\quad V_{pp\sigma}(r)=V_{pp\sigma}^{0}\ \text{exp}\big{(}-(r-d_{0})/r_{0}\big{)},\end{gathered}$ (5)
@@
Figure A1: (a) Top view of MATBG with a twist angle of $6.01^{\circ}$ ($n=5$). The lattice vectors $\bm{L}_{1}$, $\bm{L}_{2}$ and the moiré Wigner-Seitz cell are marked. (b) Triangular loops used for computing the valley charge. (c) Brillouin zones of the rotated top (blue) and bottom (orange) monolayers, and the hexagonal Brillouin zones of the coupled layers (black), drawn in an extended scheme. The $K(K^{\prime})$ points of the original graphenes are depicted. (d) Moiré Brillouin zone of MATBG, with the high-symmetry $\Gamma_{M}$, $M_{M}$, $K_{M}$, $K^{\prime}_{M}$ points and the reciprocal lattice vectors $\bm{g}_{1}$, $\bm{g}_{2}$ labeled.
Appendix B Mean-field decoupling and many-body projection
@@
$\begin{split}&{H}=H_{\text{TB}}+{H}_{\text{int}}\\ &H_{\text{TB}}=\sum_{\bm{r}\bm{r^{\prime}}s}t(\bm{r}-\bm{r^{\prime}})c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}\\ &{H}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r^{\prime}}s^{\prime}}c_{\bm{r^{\prime}}s^{\prime}}-\frac{1}{2}\bigg{)}+\frac{1}{2}\sum_{\bm{r}s}U\bigg{(}c^{\dagger}_{\bm{r}s}c_{\bm{r}s}-\frac{1}{2}\bigg{)}\bigg{(}c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}-\frac{1}{2}\bigg{)},\end{split}$ (6)
@@
$-\sum_{\bm{r}\neq\bm{r^{\prime}},s}V(\bm{r}-\bm{r})\langle c^{\dagger}_{\bm{r^{\prime}}s}c_{\bm{r}s}\rangle_{0}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\sum_{\bm{r}s}U\bigg{(}\langle c^{\dagger}_{\bm{r}\bar{s}}c_{\bm{r}\bar{s}}\rangle_{0}-\frac{1}{2}\bigg{)}c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s}+\text{constant},$ (7)
@@
$c^{\dagger}_{\bm{k}\bm{\delta}s}=\frac{1}{\sqrt{N_{c}}}\sum_{\bm{R}_{\ell}}e^{i\bm{k}\cdot(\bm{\delta}+\bm{R}_{\ell})}c^{\dagger}_{\bm{\delta}+\bm{R}_{\ell},s},$ (8)
@@
$\begin{split}{H}_{\text{MF}}=&\sum_{\bm{k}ns}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\text{constant},\\ c^{\dagger}_{\bm{k}ns}=&\sum_{\bm{\delta}}\langle\bm{k}\bm{\delta}|\bm{k}n\rangle c^{\dagger}_{\bm{k}\bm{\delta}s}=\sum_{\bm{r}}\langle\bm{r}|\bm{k}n\rangle c^{\dagger}_{\bm{r}s},\end{split}$ (9)
where $\langle\bm{k}\bm{\delta}|\bm{k}n\rangle$ and $\langle\bm{r}|\bm{k}n\rangle$ are the wave function amplitudes (independent of spin) of mode $c^{\dagger}_{\bm{k}ns}$ in the Bloch and real-space basis, respectively. The conduction bands have band index $n=1,2,...$ with increasing $n$ corresponding to increasing energy, and $n=-1,-2,...$ denotes the valence bands with increasing $|n|$ corresponding to decreasing energy. In the self-consistent algorithm we used the ground state of the bare system ($H_{\text{TB}}$) as initial seed and imposed the crystallographic symmetries and spinless time-reversal symmetry at each iteration. We used a $6\times 6$ discretization of the Brillouin zone ($N_{c}=36$); results for finer momentum grids are obtained from $H_{\text{MF}}$ (Eq. 7) with the self-consistent Fock matrix, $\langle c^{\dagger}_{\bm{r}s}c_{\bm{r^{\prime}}s^{\prime}}\rangle_{0}$, of the $6\times 6$ system. The ODA algorithm [98] is used to accelerate convergence.
@@
$\begin{split}\mathcal{P}=&\sum_{I}|\Psi_{I}\rangle\big{\langle}\Psi_{I}|,\\ |\Psi_{I}\rangle=&\prod_{\bm{k}ns\in I}c^{\dagger}_{\bm{k}ns}\prod_{\bm{q}s^{\prime},m<{-n_{B}}/{2}}c^{\dagger}_{\bm{q}ms^{\prime}}|0\rangle,\end{split}$ (10)
@@
$+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})c^{\dagger}_{\alpha}c^{\dagger}_{\beta}\prod_{j}c^{\dagger}_{j}\big{|}0\big{\rangle}+\text{others},$ (11)
@@
$\mathcal{P}V\mathcal{P}=\frac{1}{2}\sum_{\alpha\beta\gamma\delta}V_{\alpha\beta\gamma\delta}c^{\dagger}_{\alpha}c^{\dagger}_{\beta}c_{\gamma}c_{\delta}+\sum_{\alpha\beta}\sum_{i}(V_{\alpha ii\beta}-V_{i\alpha i\beta})c^{\dagger}_{\alpha}c_{\beta}+\frac{1}{2}\sum_{ik}(V_{ikki}-V_{ikik})$ (12)
Using Eq. (11) one can check that the expression in Eq. (12) has the same action as $V$ if we restrict to initial and final states contained in $\mathcal{P}$. One now sees that, crucially, the occupied modes that have been integrated out induce new terms which are precisely the Hartree and Fock potentials acting on the central bands. The constant piece is the energy of the projected-out Fermi sea.
@@
$\mathcal{P}O\mathcal{P}=\sum_{\alpha\beta}O_{\alpha\beta}c^{\dagger}_{\alpha}c_{\beta}+\sum_{i}O_{ii}.$ (13)
@@
${\mathcal{H}}=\mathcal{P}{H}\mathcal{P}=$ ${\sum_{\bm{k}ns}}^{\prime}\varepsilon^{\text{MF}}_{n}(\bm{k})c^{\dagger}_{\bm{k}ns}c_{\bm{k}ns}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\ d^{\dagger}_{\bm{r}s}d^{\dagger}_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r^{\prime}}s^{\prime}}d_{\bm{r}s}+\sum_{\bm{r}}U\ d^{\dagger}_{\bm{r}\uparrow}d^{\dagger}_{\bm{r}\downarrow}d_{\bm{r}\downarrow}d_{\bm{r}\uparrow}$
@@
$d^{\dagger}_{\bm{r}s}={\sum_{\bm{k}n}}^{\prime}\langle\bm{k}n|\bm{r}\rangle c^{\dagger}_{\bm{k}ns},\quad\quad\quad\rho_{0\bm{r}}={\sum_{\bm{k},n<0}}^{\prime}|\langle\bm{r}|\bm{k}n\rangle|^{2}.$ (15)
@@
$\begin{split}&{\mathcal{H}}={\mathcal{H}}_{0}+{\mathcal{H}}_{\text{int}}+\text{constant},\\ &{\mathcal{H}}_{0}={\sum_{\bm{k}nms}}^{\prime}{h}(\bm{k})_{nm}c^{\dagger}_{\bm{k}ns}c_{\bm{k}ms},\\ &{\mathcal{H}}_{\text{int}}=\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}},ss^{\prime}}V(\bm{r}-\bm{r^{\prime}})\delta\rho_{\bm{r}s}\delta\rho_{\bm{r^{\prime}}s^{\prime}}+\sum_{\bm{r}}U\delta\rho_{\bm{r}\uparrow}\delta\rho_{\bm{r}\downarrow}.\end{split}$ (16)
@@
$\begin{gathered}h(\bm{k})_{nm}=\varepsilon^{\text{MF}}_{n}(\bm{k})\delta_{nm}+\frac{1}{2}\sum_{\bm{r}\neq\bm{r^{\prime}}}V(\bm{r}-\bm{r^{\prime}})\langle\bm{k}n|\bm{r}\rangle\langle\bm{r^{\prime}}|\bm{k}m\rangle\bigg{(}\sideset{\,}{{}^{\prime}}{\sum}_{\bm{q},\ell<0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle-\sideset{}{{}^{\prime}}{\sum}_{\bm{q},\ell>0}\langle\bm{r}|\bm{q}\ell\rangle\langle\bm{q}\ell|\bm{r^{\prime}}\rangle\bigg{)},\\ \delta\rho_{\bm{r}s}=\rho_{\bm{r}s}-\rho_{0\bm{r}},\quad\quad\quad\quad\quad\rho_{\bm{r}s}=d^{\dagger}_{\bm{r}s}d_{\bm{r}s}.\end{gathered}$ (17)
This is the form of the effective models studied in the main text.
Appendix C Vortex Chern basis of the flat bands